OpenStudy (anonymous):

critical points: 6x−1−5y2=0 10y2−10xy=0

OpenStudy (turingtest):

I was right on my method once I remembered it. I don't blame you for doubting me though.

OpenStudy (anonymous):

hi there.. i didnt understand what you did in the other topic.. so my question is re-formulated :)

OpenStudy (anonymous):

could you please explain it more detailed so i can understand?

OpenStudy (anonymous):

but what i did was that wrong? 10y(y-x) = 0 so y must be zero

OpenStudy (turingtest):

yes that is ONE solution, but because this system is non-linear you will get three, watch:

OpenStudy (anonymous):

oke.. i will wait for your explanation :)

OpenStudy (turingtest):

sure 6x−1−5y2=0 (1) 10y2−10xy=0 (2) solve (1) for x: 6x=5y^2+1 x=(5y^2+1)/6 sub into (2): 10y^2-(10/6)(5y^2+1)y=0 10y^2-25/3y^3-5y/3=0 factor out (-5y/3): (-5y/3)(5y^2-6y+1)=0 (-5y/3)(5y-1)(y-1)=0 y={0,1/5,1} now do what you did for each y separately: for y=0: 6x-1=0->x=1/6 for y=1/5: 6x-1-5(1/5)^2=6x-1-1/5=0->x=1/5 for y=1 6x-1-5=0-->x=1 so the critical points are (0,1/6) (1/5,1/5) (1,1)

OpenStudy (turingtest):

It's always possible I made an algebraic mistake somewhere, so please check my work, but the method is kosher.

OpenStudy (anonymous):

wooh oke.. i'm gonna read it now :)

OpenStudy (anonymous):

ahhh i understand it now :) thank you for your time and explanation <3

OpenStudy (turingtest):

thanks for refreshing my memory on this stuff :)

OpenStudy (anonymous):

thank you :P